OpenStudy (anonymous):

Guys I need help please. The smallest angle in a triangle is one third of the largest angle. The third angle is ten degrees more than the largest angle. Find the measure of all three angles

jimthompson5910 (jim_thompson5910):

Let x = smallest angle y = medium sized angle z = largest angle

jimthompson5910 (jim_thompson5910):

"The smallest angle in a triangle is one third of the largest angle" means that x = (1/3)*z 3x = z z = 3x

jimthompson5910 (jim_thompson5910):

and this is a complete contradiction "The third angle is ten degrees more than the largest angle" How can you have an angle larger than the largest angle?

OpenStudy (anonymous):

No

jimthompson5910 (jim_thompson5910):

if it was something like "the middle angle is ten degrees LESS than the largest angle", then it would make more sense

OpenStudy (anonymous):

It is smallest sorry

jimthompson5910 (jim_thompson5910):

if it's "The third angle is ten degrees more than the SMALLEST angle", then y = x+10

jimthompson5910 (jim_thompson5910):

the three angles in the triangle always add to 180 degrees angle1+angle2+angle3 = 180 x+y+z = 180 x+(x+10)+(3x) = 180 ... plug in y = x+10 and z = 3x x+x+10+3x = 180 x = ???

OpenStudy (anonymous):

3?

jimthompson5910 (jim_thompson5910):

x+x+10+3x = 180 5x+10 = 180 ... ... ... x = ???

OpenStudy (anonymous):

18

jimthompson5910 (jim_thompson5910):

5x+10 = 180 5x = 180-10 5x = 170 so x must be ?

jimthompson5910 (jim_thompson5910):

5x means "5 times x"

OpenStudy (anonymous):

850

jimthompson5910 (jim_thompson5910):

5 * _______ = 170 what goes in the blank?

OpenStudy (anonymous):

34

jimthompson5910 (jim_thompson5910):

good, you find this by dividing both sides by 5

jimthompson5910 (jim_thompson5910):

then take this x value and plug it into the equations below y = x+10 z = 3x

OpenStudy (anonymous):

Ok